OpenStudy (anonymous):

The width of a rectangle is 12 units. Can the perimeter x of the rectangle can be 60 units when its length y is 18 units? No, the rectangle cannot have x = 60 and y = 18 because x + y is less than 60. No, the rectangle cannot have x = 60 and y = 18 because x + y is less than 24. Yes, the rectangle can have x = 60 and y = 18 because x = 12 + 2y. Yes, the rectangle can have x = 60 and y = 18 because x = 24 + 2y.

OpenStudy (texaschic101):

Perimeter of rectangle formula is : P = 2L + 2W...however, they have x as the perimeter and y(the length) as 18....so your equation using these letters is : X = 2y + 2W so we have : X = (perimeter) y (length) = 18 W (width) = 12 so lets sub and find X (the perimeter).. X = 2y + 2W X = 2(18) + 2(12) X = 36 + 24 X = 60 yes..the rectangle can have X = 60 and y = 18 because X = 24 + 2y check.. X = 24 + 2y 60 = 24 + 2(18) 60 = 24 + 36 60 = 60 (correct)

OpenStudy (anonymous):

The answer would be D